A tube is spewing sand at a rate of one cubic meter per second. It generates a cone-shaped pile of material. The radius of the circle at the base of the cone is equal to its height. When the sand pile reaches 2 meters in height, how fast does it rise?


1
Expert's answer
2021-10-29T03:02:17-0400

The volume of cone is

"V=\dfrac{1}{3}\pi r^2h"

If "h=r," then


"V=\dfrac{1}{3}\pi h^3"

Differentiate both sides with respect to "t"


"\dfrac{dV}{dt}=\dfrac{d}{dt}(\dfrac{1}{3}\pi h^3)"

Use the Chain Rule


"\dfrac{dV}{dt}=\pi h^2\dfrac{dh}{dt}"

Solve for


"\dfrac{dh}{dt}=\dfrac{\dfrac{dV}{dt}}{\pi h^2}"

Given "\dfrac{dV}{dt}=1m^3/s, h=2m."


"\dfrac{dh}{dt}=\dfrac{1m^3/s}{\pi (2m)^2}=\dfrac{1}{4\pi}\ m/s\approx0.080\ m/s"

When the sand pile reaches 2 meters in height, it raises at speed of "\dfrac{1}{4\pi}\ m/s\approx0.080\ m/s."